Form the pair of linear equations for the following problems and find their solution by substitution method.
 
The larger of two supplementary angles exceeds the smaller by 18 degrees. Find them.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the larger angle by x and smaller angle be y .(in degrees)
We know that the sum of two supplementary pair of angles is always 180 degrees .
According to the question
x + y = 180……………. (1)
x – y = 18……………..(2)
From (1),
we get x = 180 – y …………. (3)
Substituting (3) in (2)
we get
180 – y – y =18
162 = 2y
y = 81 ………….. (4)
Using the value of y in (3),
we get x = 180 – 81 = 99
Hence, the angles are 99 and 81 .

Thanks.